OpenStudy (anonymous):

Which ratio is Proportional to 3/5 ?

OpenStudy (whpalmer4):

Just about all of them, in general :-) To compare ratios or fractions, cross-multiply, multiplying the numerator of one by the denominator of the other, and vice versa. If the two products are not equal, the fractions are not equal. For example: \(\dfrac{2}{3} = \dfrac{3}{4}\) only if \(2*4\) = \(3*3\). As it does not, the fractions are not equal. \(\dfrac{2}{3} = \dfrac{4}{6}\) only if \(2*6\) = \(3*4\). They are, so the fractions are equal.

OpenStudy (anonymous):

Wow Thank you man

OpenStudy (whpalmer4):

Hey, you changed the question on me! Answer is still valid, but I would strike the opening comment.